OpenStudy (luigi0210):

IBP:

OpenStudy (luigi0210):

\[\LARGE \int t~sec^2(2t)~dt\]

zepdrix (zepdrix):

\[\Large\rm \int\limits \sec^2(2t)dt\]

zepdrix (zepdrix):

Like do you know how to deal with the inner 2, without making an extra substitution?

zepdrix (zepdrix):

blah my message didn't post, it said: "quick question: can you do this integral in your head?"

OpenStudy (luigi0210):

Would it be \(\Large \frac{1}{2}tan(2t)+C\)?

zepdrix (zepdrix):

Yes, good. We end up with a 1/2 because of the 2.

zepdrix (zepdrix):

So that takes care of our dv and v.\[\Large\rm dv=\sec^2(2t), \qquad\qquad v=\frac{1}{2}\tan(2t)\]

zepdrix (zepdrix):

Err I should be a little more accurate and say:\[\Large\rm dv=\sec^2(2t)dt\]

zepdrix (zepdrix):

\[\Large\rm u=t, \qquad\qquad \quad dv=\sec^2(2t)dt\]\[\Large\rm du=dt,\qquad\qquad v=\frac{1}{2}\tan(2t)\]

zepdrix (zepdrix):

Understand why I chose to for the u?

OpenStudy (luigi0210):

Yup, yup~

zepdrix (zepdrix):

So where ya stuck? >:U Just do it!!

OpenStudy (luigi0210):

Just to clarify, after all that it would end up as \[\LARGE \frac{t}{2}tan(2t) -\frac{1}{2} \int tan(2t)dt\]?

zepdrix (zepdrix):

Mmmmmm ya looks good.

zepdrix (zepdrix):

Remember how to integrate tangent? That one's not too bad.

OpenStudy (luigi0210):

Uhm, \(\LARGE \int~tanx~dx=ln|sec~x|+C\)?

zepdrix (zepdrix):

Mmm ya that sounds right.

OpenStudy (luigi0210):

Okay, thank you zepdrix :D